# A Construction of Infinite Families of Self-Orthogonal Quasi-Cyclic Codes Using Constituent Codes

### v1

@@ 5. Conclusion
In general, describing the parameters of infinity families of codes is not an easy task. In this paper, we have presented families of QC codes over $\mathbb{F}_{q}$, making explicit their dimension and a lower bound to their minimum distances. In addition, based on their constituent code decomposition, we also explored their dualities according to Euclidean and Hermitian inner products. We have also addressed when this lower bound to minimum distance is greater than the square-root-like bound, which there are only a few constructions in the literature achieving this goal. Finally, some quantum codes have been reported, which are good ones in comparison with those ones depicted at the Code Table [15].

### v2

@@ 5. Conclusion
In general, describing the parameters of infinity families of codes is not an easy task. In this paper, we have presented families of QC codes over $\mathbb{F}_{q}$, making explicit their dimension and a lower bound to their minimum distances. In addition, based on their constituent code decomposition, we also explored their dualities according to Euclidean and Hermitian inner products. We have also addressed when this lower bound to minimum distance is greater than the square-root-like bound, which there are only a few constructions in the literature achieving this goal. Finally, some quantum codes have been reported, which are comparable with those ones depicted on the Code Table [table].